OpenStudy (anonymous):

if x/2y=6/7 then the value of x-y/x+y + 14/19 is wat?

OpenStudy (anonymous):

so you can say that: \[x = \frac{12y}{7}\]

hero (hero):

Yup, you solved it

OpenStudy (anonymous):

you could plug this into the equation: \[\frac{\frac{5y}{7}}{\frac{19y}{7}}\]

OpenStudy (anonymous):

whoever wrote this problem is the devil

OpenStudy (anonymous):

the answer is 1

OpenStudy (anonymous):

the 7s and the ys cancel

hero (hero):

lol

OpenStudy (anonymous):

you get: \[\frac{5}{19} + \frac{14}{19} = 1\]

hero (hero):

Good job

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thanks.....

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

did my answer seem to jump to conclusions?

OpenStudy (anonymous):

is there anything you want explained more?

OpenStudy (anonymous):

na.... I understood the reply.... tats whts required.... ur answer is precise to the point....

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

in fact I'm new here.... I jus wanted to check out how r things here.... and by the looks of it.... I like this site....